What is the big o cost of this method?

int count = 0;
int i = 1;
while(i < n){
    for (j = 1; j < n*n; j *= n) {
          count++; 
    }
    i *= 2;
}
System.out.println(count); 
Answer #1
Number of values for i = log(n)
Number of values for j = logn(n^2) = 2

So, Time complexity 
= O(log(n)*2)
= O(log(n))

O(log(n))